OpenStudy (anonymous):

2x-10/x^2-3x-10 multiplied by x^2-4/8x-16

OpenStudy (anonymous):

Start by factoring

OpenStudy (solomonzelman):

factor completely, cancel terms out, multiply

OpenStudy (anonymous):

but the 8x-16 part is getting me confused what would it cancel with

OpenStudy (solomonzelman):

factor out of 8

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{2x-10}{x^2-3x-10} \times \frac{x^2-4}{8x-16} }\) \(\LARGE\color{blue}{ \frac{2(x-5)}{(x-5)(x+2)} \times \frac{(x-2)(x+2)}{8(x-4)} }\)